OpenStudy (anonymous):

Assume that a 15-kg ball moving at 8 m/s strikes a wall perpendicularly and rebounds elastically at the same speed. What is the amount of impulse given to the wall? Show your work. I'm inexperienced... walk me through this?

OpenStudy (akashdeepdeb):

Do you know that Force is Change in momentum / time ? Because F = ma = m(v-u)/t = (mv - mu)/t = change in momentum/time!! :D Thus Impulse = F*t = change in momentum * time/time = change in momentum!! :D So Impulse is change in momentum!  m1u1 + m2u2 = m1v1 + m2v2 [Law of conservation of momentum] Now the wall is still so u2,v2 = 0 or m1u1 = m1v1 Now we know that as it is elastic! So even kinetic energy is conserved!! 1/2m1u1^2 = 1/2m1v1^2 [u2,v2 = 0] u1 = v1 [Just direction should be opposite] m1u1 = 15 * 8 = 120kg.m/s m1v1 = -15 * 8 = -120kg.m/s Now we have to calculate impulse = Change in momentum! = m1v1 - m1u1 = -120 - 120 = -240kg.m/s!! :D Get it? :)

OpenStudy (akashdeepdeb):

btw in the drawing it is law of conservation of momentum! And law of conservation of kinetic energy as the collision is elastic!! :D

OpenStudy (anonymous):

O_O Lol, thank you, much appreciated.

OpenStudy (akashdeepdeb):

:)